# Optimal Baseline Corrections for Off-Policy Contextual Bandits

### v1

@@ 6. Conclusion and Future Work
In this work, we have proposed to unify different off-policy estimators as equivalent additive baseline corrections. We look at off-policy evaluation and learning settings and propose baseline corrections that minimize the variance in the estimated policy value and the empirical gradient of the off-policy learning objective. Extensive experimental comparisons on a synthetic benchmark with realistic settings show that our proposed methods improve performance in the OPE and OPL tasks.
We believe our work represents a significant step forward in the understanding and use of off-policy estimation methods (for both evaluation and learning use-cases), since we show that the prevalent SNIPS estimator can be improved upon with essentially no cost, as our proposed method is parameter-free and – in contrast with SNIPS – it retains the unbiasedness that comes with IPS.
@@ Appendix A Appendix: Off-policy Estimator Variance
In this appendix, we report additional results from the experimental section (Section 4 from the main paper), answering RQ 4. Specifically, we look the the empirical variance of various offline estimators for the task of off-policy evaluation. The mean squared error (MSE) of different offline estimators are reported in Figure 4 of the main paper. In this appendix, we report the empirical variance of various offline estimators in Figure 5.
From the figure, it is clear that our proposed $\beta$-IPS estimator with estimator variance minimizing $\beta$ value (Eq. 39 in the main paper) results in the lowest empirical variance in most of the cases. It is interesting to note that when the logged data is limited ($N<10^{3}$), sometimes the SNIPS estimator has lower estimator variance. We suspect that the reason could be a bias in the estimate of the variance-optimal $\beta$ estimate (Eq. 38 in the main paper), when the dataset size is small, given that it is a ratio estimate of expectations. For practical settings, i.e. when $N>10^{3}$, the proposed estimator $\beta$-IPS results in a minimum sample variance, thereby empirically validating the effectiveness of our proposed $\beta$-IPS estimator for the task of OPE.

### v2

@@ 6. Conclusion and Future Work
In this work, we have proposed to unify different off-policy estimators as equivalent additive baseline corrections. We look at off-policy evaluation and learning settings and propose baseline corrections that minimize the variance in the estimated policy value and the empirical gradient of the off-policy learning objective. Extensive experimental comparisons on a synthetic benchmark with realistic settings show that our proposed methods improve performance in the off-policy estimation (OPE) and off-policy learning (OPL) tasks.
We believe our work represents a significant step forward in the understanding and use of off-policy estimation methods (for both evaluation and learning use-cases), since we show that the prevalent SNIPS estimator can be improved upon with essentially no cost, as our proposed method is parameter-free and — in contrast with SNIPS — it retains the unbiasedness that comes with IPS.
@@ Appendix A Appendix: Off-policy Estimator Variance
In this appendix, we report additional results from the experimental section (Section 4 from the main paper), answering RQ4. Specifically, we look the the empirical variance of various offline estimators for the task of off-policy evaluation. The mean squared error (MSE) of different offline estimators are reported in Figure 4. In this appendix, we report the empirical variance of various offline estimators in Figure 5.
From the figure, it is clear that our proposed $\beta$-IPS estimator with estimator variance minimizing $\beta$ value (Eq. 38) results in the lowest empirical variance in most of the cases. It is interesting to note that when the logged data is limited ($N<10^{3}$), sometimes the SNIPS estimator has lower estimator variance. We suspect that the reason could be a bias in the estimate of the variance-optimal $\beta$ estimate (Eq. 38), when the dataset size is small, given that it is a ratio estimate of expectations. For practical settings, i.e., when $N>10^{3}$, the proposed estimator $\beta$-IPS results in a minimum sample variance, thereby empirically validating the effectiveness of our proposed $\beta$-IPS estimator for the task of OPE.
Acknowledgements
The authors would like to thank Adith Swaminathan, Thorsten Joachims, Ben London, and Philipp Hager for valuable discussions and feedback on early drafts of this manuscript.
This research was supported by Huawei Finland, the Dutch Research Council (NWO), under project numbers VI.Veni.222.269, 024.004.022, NWA.1389.20.183, and KICH3.LTP.20.006, and the European Union’s Horizon Europe program under grant agreement No 101070212. All content represents the opinion of the authors, which is not necessarily shared or endorsed by their respective employers and/or sponsors.